OpenStudy (anonymous):

solve the equation using the quadratic formula: x^2+4=4x

OpenStudy (anonymous):

\[ x^2+4 =4 x\\ x^2-4 x + 4 =0\\ (x-2)^2=0\\ x=2 \]

OpenStudy (anonymous):

Do it with the quadratic formula and confirm the answer above.

OpenStudy (anonymous):

I need help with setting up the quadratic formula i can't get it set up right.

OpenStudy (anonymous):

\[ x^2-4 x + 4=0\\ a=1; b=-4; c=4\\ \Delta = b^2-4 a c= 16- 4(1)(4)=0\\ x=\frac { -b\pm \sqrt \Delta}{2 a}=\frac { 4\pm 0}{2}=2 \]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

YW